# On Creating A Brain-To-Text Decoder

### v1

@@ 1 Introduction
Previous approach have been limited on classifying information need Michalkova et al. (2022) , mental work load Kingphai and Moshfeghi (2021, 2023), a reading task Hollenstein et al. (2021) or an imagined categoryNieto et al. (2022) to name a few . To our knowledge, this study is the first attempt to use the raw EEG signals for the purpose of identifying the exact words and generate sentences derived from the human brain. Our approach is significant because EEG is a real-time, less expensive, and easier-to-use technique compared to fMRI. By decoding the human brain using raw EEG signals, we aim to speed up the process of understanding the human brain using a faster and more efficient method. We also aim in our knowledge to be the first that utilise raw EEG signals as the input to our model to perform the decoding.
@@ 2 Related work
Recent works by Kostas et al. (2021a) and Partovi et al. (2023) have explored the potential to produce more potent and generalized EEG representations. BENDR, introduced by Kostas et al. (2021a), employs transformer architectures and contrastive self-supervised learning to enhance the generation of general EEG embeddings. This approach aims to capture intricate temporal dependencies and patterns within the data. On the other hand, Partovi et al. (2023) presents a self-supervised learning framework focused on creating task-agnostic EEG embeddings, enhancing adaptability for various downstream applications.

### v2

# On Text Generation Using Brain Data

@@ 1 Introduction
Previous approach have been limited on classifying information need Michalkova et al. (2022) , mental work load Kingphai and Moshfeghi (2021b, 2023), a reading task Hollenstein et al. (2021) or an imagined categoryNieto et al. (2022) to name a few . To our knowledge, this study is the first attempt to use the raw EEG signals for the purpose of identifying the exact words and generate sentences derived from the human brain. Our approach is significant because EEG is a real-time, less expensive, and easier-to-use technique compared to fMRI. By decoding the human brain using raw EEG signals, we aim to speed up the process of understanding the human brain using a faster and more efficient method. We also aim in our knowledge to be the first that utilise raw EEG signals as the input to our model to perform the decoding.
@@ 2 Related work
Recent interdisciplinary research has explored the potential of neuroscience to advance information retrieval systems, conceptualized as the emerging field of NeuraSearch Moshfeghi (2021). Findings from research Mueller et al. (2008); Müller-Putz et al. (2015); Gwizdka et al. (2019); Martínez-Castaño et al. (2022) suggest that NeuraSearch seeks to enhance information accessibility and utilization by employing neurophysiological signals to deliver implicit relevance feedback, thereby augmenting user interaction in the absence of explicit inputs. Studies such as Moshfeghi et al. (2013) have built the foundation of NeuraSearch. For example, in Moshfeghi et al. (2016), Moshfeghi et al utilized fMRI to explore the cerebral processes that underlie information needs during Q/A tasks, thereby mapping neural activities associated with various states of knowledge. This body of research highlights the importance of neurological insights in the customization of user experiences in information retrieval, as further validated by EEG studies like Moshfeghi et al. (2019), which identified EEG patterns associated with different stages of information need recognition.
The study by Michalkova et al. (2024) explored the enhancement of queries through the utilization of brain signals, highlighting that models driven by electroencephalography (EEG) can improve the relevance of search outcomes. This finding emphasizes the practical advantages of combining traditional query strategies with neurophysiological information to optimize search processes.
NeuraSearch prioritizes the evaluation of cognitive workload through EEG analysis, with research emphasizing its efficacy in monitoring cognitive states across various contexts Kingphai and Moshfeghi (2021a). Investigations by Kingphai and Moshfeghi (2021b, 2023) have achieved progress in the areas of emotion recognition and EEG data classification, highlighting the criticality of methodological accuracy and the necessity for standardized practices. Prominent neuroimaging modalities, including MEG and fMRI, are integral to NeuraSearch’s explorations Kauppi et al. (2015). Their application, as illustrated in Lamprou et al. (2022), explored semantic comprehension and demonstrated how neuroscience can proficiently assess cognitive processing and decision-making.
The convergence of neuroscience with information retrieval elucidates unprecedented possibilities for the development of adaptive, user-centered systems. By proficiently interpreting users’ implicit feedback through neurophysiological signals, as evidenced in several pivotal studies Allegretti et al. (2015); Jacucci et al. (2019), it facilitates the creation of more nuanced interfaces. This ongoing synthesis aims to address complex user information requirements and adapt to cognitive challenges inherent in the process of information retrieval.
Moreover recent works by Kostas et al. (2021a) and Partovi et al. (2023) have explored the potential to produce more potent and generalized EEG representations. BENDR, introduced by Kostas et al. (2021a), employs transformer architectures and contrastive self-supervised learning to enhance the generation of general EEG embeddings. This approach aims to capture intricate temporal dependencies and patterns within the data. On the other hand, Partovi et al. (2023) presents a self-supervised learning framework focused on creating task-agnostic EEG embeddings, enhancing adaptability for various downstream applications.